OpenStudy (anonymous):

The minimum value of the expression 2/(15 + 2sinx - 2sqrt3 cosx)

OpenStudy (anonymous):

have you differentiated it

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

diff the denominator!

OpenStudy (anonymous):

well to find the min you have to differentiate it once; to find x; then differentiate it again to determine if its min. or max. why not you try and differentiate it ?>

OpenStudy (experimentx):

maximize this value (2sinx - 2sqrt3 cosx)

OpenStudy (anonymous):

\[\large\text{Keep it in your mind}\] \[-\sqrt{a^2+b^2}\le a\sin x+b\cos x\le\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

wat is this!

OpenStudy (anonymous):

as Exper said u need to maximize (2sinx - 2sqrt3 cosx) is that right?

OpenStudy (experimentx):

this is pretty old trick ... and very useful one!!

OpenStudy (anonymous):

yeah... @Yahoo!

OpenStudy (anonymous):

sin is max at 1

OpenStudy (anonymous):

so i have to subs that value here

OpenStudy (anonymous):

r u there guys!!

OpenStudy (anonymous):

yes yahoo u want to maximize (2sinx - 2sqrt3 cosx) according to the formula i gave u...here \(a=2\) and \(b=-2\sqrt{3}\)

OpenStudy (experimentx):